OpenStudy (anonymous):

A rectangular porch has a perimeter of 49 ft and a width of 6.5 ft find the area of the porch in square feet

OpenStudy (anonymous):

(49/2-6.5) 6.5 = ?

OpenStudy (anonymous):

Length = (49 - 2*6.5)/2 = 36/2 = 18. Area = length * width = 18 * 6.5 = ?

OpenStudy (anonymous):

you know the width, so you need the length. since the perimeter is 49 and \[2\times l +2\times 6.5=49\] you can find the length. then multiply to find the area